Building Lot A requires $50,000 down plus an additional $100 per square foot. Building Lot B requires $60,000 down plus an additional $50 per square foot. What size lot would make the two lots the same price?

Answer by Boreal(15235)

50000+100x=60000+50x, where x is the number of square feet.
50x=10000
x=200 sq ft. ANSWER
50000+10000=60000+10000